OpenStudy (superhelp101):

An equation of a circle is x2 + y2 + 10x + 6y + 18 = 0. Show all your work in determining the center and radius of this circle. In complete sentences, explain the procedure used.

OpenStudy (superhelp101):

@jim_thompson5910

OpenStudy (superhelp101):

@aum

jimthompson5910 (jim_thompson5910):

we have x^2 and 10x what else is needed to make it a perfect square?

jimthompson5910 (jim_thompson5910):

x^2 + 10x + ______

OpenStudy (superhelp101):

10x ?

OpenStudy (superhelp101):

why 10x ?

jimthompson5910 (jim_thompson5910):

it's in the original equation

jimthompson5910 (jim_thompson5910):

I'm pulling all the x terms together

OpenStudy (superhelp101):

oh okay, so we are taking x^2+10x. Would we bring y^2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

don't worry about the y terms right now

OpenStudy (superhelp101):

oh okay

jimthompson5910 (jim_thompson5910):

what goes in the blank x^2 + 10x + ______ to make that a perfect square trinomial

OpenStudy (superhelp101):

+18

jimthompson5910 (jim_thompson5910):

try again

OpenStudy (superhelp101):

+0

OpenStudy (superhelp101):

is it from the equation?

jimthompson5910 (jim_thompson5910):

Focus on 10x what's half of 10?

OpenStudy (superhelp101):

5. But can't it also be 25 ?

jimthompson5910 (jim_thompson5910):

then square 5 to get 25 so x^2 + 10x + 25 is a perfect square it factors to (x+5)^2

OpenStudy (superhelp101):

yes

jimthompson5910 (jim_thompson5910):

So add 25 to both sides of the original equation to get x^2 + y^2 + 10x + 6y + 18 = 0 x^2 + y^2 + 10x + 6y + 18 + 25 = 0+25 (x^2+ 10x + 25) + y^2 + 6y + 18 = 25 (x+5)^2 + y^2 + 6y + 18 = 25

jimthompson5910 (jim_thompson5910):

you must do the same for the y terms as well

OpenStudy (superhelp101):

oh that makes sense so far.

jimthompson5910 (jim_thompson5910):

y^2 + 6y + _____

OpenStudy (superhelp101):

9

OpenStudy (superhelp101):

I just cut 6y in half and squared it.

jimthompson5910 (jim_thompson5910):

add 9 to both sides of (x+5)^2 + y^2 + 6y + 18 = 25

jimthompson5910 (jim_thompson5910):

and complete the square for y to factor

OpenStudy (superhelp101):

(x+5)^2 + y^2 + 6y + 18+9 = 25+9

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (superhelp101):

okay so: (x+5)^2 + y^2 + 6y +27=35 x^2+10x+25+ y^2 + 6y +27=35 x^2+ y^2+10x+ 6y +52=35 Did i do it right?

jimthompson5910 (jim_thompson5910):

don't combine 18+9 to get 27

jimthompson5910 (jim_thompson5910):

move that +9 over with the y^2+6y

jimthompson5910 (jim_thompson5910):

you're going to factor y^2+6y+9 to get (y+3)^2

jimthompson5910 (jim_thompson5910):

also, 25+9 = 34, not 35

OpenStudy (superhelp101):

wait I am so confused, can you show me :/

jimthompson5910 (jim_thompson5910):

(x+5)^2 + y^2 + 6y + 18+9 = 25+9 (x+5)^2 + y^2 + 6y+9 + 18 = 25+9 (x+5)^2 + (y^2 + 6y+9) + 18 = 25+9 (x+5)^2 + (y+3)^2 + 18 = 34

jimthompson5910 (jim_thompson5910):

you still have more work to do though

OpenStudy (superhelp101):

okay

jimthompson5910 (jim_thompson5910):

You have to get the equation into the form \[\Large (x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (superhelp101):

so subtract 18 from both sides, and square it?

jimthompson5910 (jim_thompson5910):

yes to the first part, no to the second

jimthompson5910 (jim_thompson5910):

(x+5)^2 + (y+3)^2 + 18 = 34 (x+5)^2 + (y+3)^2 + 18-18 = 34-18 (x+5)^2 + (y+3)^2 = 16

jimthompson5910 (jim_thompson5910):

what's next

OpenStudy (superhelp101):

I think take the square root of 16 ?

jimthompson5910 (jim_thompson5910):

yep sqrt(16) = 4

jimthompson5910 (jim_thompson5910):

so 4^2 = 16

jimthompson5910 (jim_thompson5910):

r^2 = 4^2 = 16 therefore, r = 4 the radius is 4

OpenStudy (superhelp101):

ohh makes sense.

jimthompson5910 (jim_thompson5910):

what's the center?

OpenStudy (superhelp101):

umm.(-5,-3)...but I have no clue I got that

jimthompson5910 (jim_thompson5910):

(x+5)^2 + (y+3)^2 = 16 (x+5)^2 + (y+3)^2 = 4^2 compare to \[\Large (x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (superhelp101):

wait so it's (-5,-3) ?

OpenStudy (superhelp101):

do i set them to zero?

jimthompson5910 (jim_thompson5910):

how do you get (x+5)^2 into the form (x-h)^2

OpenStudy (superhelp101):

i don't know :(

jimthompson5910 (jim_thompson5910):

notice we have x PLUS 5 and we want a MINUS h

jimthompson5910 (jim_thompson5910):

so x+5 = x - (-5) now that's in the form x-h

jimthompson5910 (jim_thompson5910):

h = -5

OpenStudy (superhelp101):

so we find y-k the same way ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (superhelp101):

oh okay

jimthompson5910 (jim_thompson5910):

k = ???

OpenStudy (superhelp101):

-3

OpenStudy (superhelp101):

y+3=y-(-3) y=-3 Right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

k = -3

jimthompson5910 (jim_thompson5910):

h = -5 k = -3

jimthompson5910 (jim_thompson5910):

(h,k) is the center

OpenStudy (superhelp101):

yes, thank you for your time, I'm sorry if I took up ur time :(.

jimthompson5910 (jim_thompson5910):

you didn't and you're welcome